OpenStudy (animalover02):

I WILL METAL AND FAN!!!!!!!!!!!!! which set of ordered pairs contains only points that are on the graph of the function y=12-3x? A. (-3 , -27) , (0 , 0 ) , (6 , 54) B. (-18 , 10) , ( -6 , 6 ) , (18 , -2) C. (-5 , 27) , (-1 , 15) , ( 8 , -12 ) D. ( -7 , -9 ) , (-4 , 0 ) , ( 2 , 18 )

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

your math teacher must hate you!

OpenStudy (animalover02):

hey:) lol

OpenStudy (misty1212):

you have to check for each one if the second number is \[12-3x\] where \(x\) is the first number so you got a lots to do

OpenStudy (misty1212):

for example, for \((-3,-27\) you have to check if \[12-3\times (-3)=-27\]

OpenStudy (animalover02):

i know how to do these but i take forever on them . my test shuts off in 2 minutes so i needed someone whos fast to do it:)

OpenStudy (alivejeremy):

Medal*

OpenStudy (misty1212):

it doesn't because \[12-3\times (-3)=12+9=21\]so NO to that one

OpenStudy (animalover02):

so A is incorrect?

OpenStudy (misty1212):

\[12-3\times 0=0?\] NO

OpenStudy (misty1212):

yeah A is a NO B is a NO

OpenStudy (animalover02):

ok:)

OpenStudy (misty1212):

\[54=12-3\times 6\]? not likely is it?

OpenStudy (skullpatrol):

A fast way would be to graph the line and plot the points

OpenStudy (animalover02):

so D?

OpenStudy (animalover02):

i wish you could metal more than 1 person but you cant

OpenStudy (misty1212):

D would make \[12-3\times (-18)=10\]

OpenStudy (misty1212):

they are trying to trick you, but no it is not true

OpenStudy (animalover02):

so its C

OpenStudy (misty1212):

it is true that \[12-3\times 10=-18\] but that is backwards

OpenStudy (skullpatrol):



OpenStudy (animalover02):

C or D?

OpenStudy (misty1212):

stop with the A, B, C, D you just gotta check each one

OpenStudy (misty1212):

so far, none have worked

OpenStudy (animalover02):

i have 15 sec. please need answer i will fan all of you please

OpenStudy (misty1212):

ii can't do it in 15 seconds sorry

OpenStudy (skullpatrol):

:(

OpenStudy (skullpatrol):

you could do it really fast with the graphing method next time

OpenStudy (skullpatrol):

you only have to check the points that are near the line